OpenStudy (anonymous):

Solve the system of equations y=2x^2-3 y=3x-1 No solution (-1/2,5), (2,-5/2) (-1/2,-5/2), (2,5) (1/2,5/2), (2,5)

OpenStudy (gorv):

3x-1=2x^2-3 2x^2-3x-2=0 2x^2-4x+x-2=0 (2x+1)(x-2)=0 x=-1/2 and x=2 x=-1/2 y=-5/2 x=2 y=5 solution (-1/2,-5/2) (2,5)

OpenStudy (gorv):

@indianwaking97

OpenStudy (anonymous):

Thanks